OpenStudy (nafisa):

What is the relationship between radius of bohr orbit and principal quantum number?

OpenStudy (osprey):

Enclosed is a jpg picture of the equation I think you want ... Also, a website you might find helpful (it's intended to be) http://perendis.webs.com

OpenStudy (nafisa):

Thanks

OpenStudy (sharmagk83):

Radius of Bohr's orbit is given by \[r = \frac{n ^{2}h^{2} \epsilon_{0} }{ \pi m e^{2}}\] where \[ \frac{h^{2} \epsilon_{0} }{ \pi m e^{2}} = constant\] say K so we can write as \[r = K n^{2}\]